# Understanding the Complexities of Responsibly Sharing NSFW Content Online

### v1

@@ Appendix B Ethics Statement
  • Confidentiality: To safeguard the privacy of Reddit users, we refrained from performing any cross-platform linking. Additionally, our analysis did not focus on user-specific patterns but instead centered on a general examination of NSFW subreddit. We did not collect any videos or images associated with the posts. To ensure the privacy of Reddit users, our study was strictly limited to text of posts and associated metadata.
  • Potential for harm: We implement a protocol to notify competent authorities to take action in case we observe any illegal content within subreddit. Our aim is not to harm Reddit users in any way. Our ethics statement ensures that our reporting focuses solely on identifying potential posts suggesting non-consensual sharing, without disclosing any Reddit user IDs or post IDs.

### v2

@@ Appendix B Ethics Statement
  • Confidentiality: To safeguard the privacy of Reddit users, we refrained from performing any cross-platform linking. Additionally, our analysis did not focus on user-specific patterns but instead centered on a general examination of NSFW subreddit. We did not collect any videos or images associated with the posts. To ensure the privacy of Reddit users, our study was strictly limited to the text of posts and associated metadata.
  • Potential for harm: We implement a protocol to notify competent authorities to take action in case we observe any illegal content within the subreddit. Our aim is not to harm Reddit users in any way. Our ethics statement ensures that our reporting focuses solely on identifying potential posts suggesting non-consensual sharing, without disclosing any Reddit user IDs or post IDs.